OpenStudy (savannah_noelle):

what is the axis of symmetry for f(x)=-2x^2+20x-42?

OpenStudy (campbell_st):

for a parabola \[y = ax^2 + bx + c\] the axis of symmetry is always \[x = \frac{-b}{2a}\] in our question b = 20 and a = -2... so just substitute and evaluate

OpenStudy (savannah_noelle):

x= -20/2(-2)

OpenStudy (campbell_st):

yep just evaluate it for the axis...

OpenStudy (savannah_noelle):

So x=5? Thank you

OpenStudy (campbell_st):

yes that's it. Glad to help.

OpenStudy (savannah_noelle):

@campbell_st Can you help me with this problem? what are the coordinates of the vertex for f(x)=2x^2+4x+9?

OpenStudy (anonymous):

dont help did black nigg

OpenStudy (savannah_noelle):

?..